OpenStudy (anonymous):

Find the following. Assume variables represent any real number. sqrt(a+4)^2=

OpenStudy (angela210793):

=|a+4| or a+4 depends is it all sqr in the 2nd power or just a+4?

OpenStudy (angela210793):

\[\sqrt{(a+4)}^2=|a+4|\] and \[(\sqrt{a+4})^2=a+4\]

OpenStudy (anonymous):

this can be done like this \[\sqrt{(a+4)^{2}}=|a+4|\] here |a+4|=a+4 if a+4>0 and hence a>-4 and |a+4|=-(a+4) if a+4<0 and hence a<-4